IF MY AGE IS 2 YEAR MY SISTER IS HALF AGE OF MY AGE IF I ' M 50 YEAR OLD WHAT WILL BE MY SISTER AGE plz tell me ans.......
Answer by ptaylor(2198)

You are 2 years old in 48 years, you will be 48+2=50 years old
Your sister is(1/2)*2=1 year old, in 48 years, she will be 48+1=49 years old

There will always be 1 year difference in your age.
Hope this helps----ptaylor